# Randomization Tests for Conditional Group Symmetry

### v1

@@ 2.2 Testing for distributional invariance
We briefly review the basic framework and properties of tests for group invariance of a probability measure. A function $f$ with domain $\mathbf{X}$ is $\mathbf{G}$-invariant if $f(gx)=f(x)$, all $x\in\mathbf{X},g\in\mathbf{G}$. This specializes to probability measures.
@@
A probability measure $P$ on $\mathbf{X}$ is $\mathbf{G}$-invariant if $g_{*}P(A)\coloneq P(g^{-1}A)=P(A)$ for all $g\in\mathbf{G}$, $A\in\mathbf{S}_{\mathbf{X}}$.
For a specified finite or compact group $\mathbf{G}$, consider the hypotheses
[…]
Following Koning and Hemerik (2024), a level-$\alpha$ test based on statistic $T$ is
$\phi_{\alpha}(X)=\mathbbm{1}\{T(X)>q_{T(GX)}(1-\alpha)\}\;,$ (2)
where $q_{T(GX)}(1-\alpha)$ is the quantile function of $T(GX)$ at $1-\alpha$, with $X\sim P$ and $G\sim\lambda$, a random element of $\mathbf{G}$ sampled from Haar measure on $\mathbf{G}$. Showing that this test has level $\alpha$ is straightforward, using the fact that $P$ is $\mathbf{G}$-invariant if and only if $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}GX$. Hence $\mathbb{E}[\phi_{\alpha}(X)]=\mathbb{E}[\phi_{\alpha}(GX)]\leq\alpha$.
@@
$p_{B}\coloneqq\frac{1+\sum_{b=1}^{B}\mathbbm{1}\{T(X_{1:n}^{(b)})\geq T(X_{1:n})\}}{1+B}$ (3)
@@
For small and finite $\mathbf{G}$, the quantile $q_{T(GX)}(1-\alpha)$ can be computed exactly by applying each element of $\mathbf{G}$ to $X$. However, in the case that $\mathbf{G}$ is large or uncountable, randomization tests may be used. Algorithm 1 shows the basic recipe. We note that Algorithm 1 differs slightly from the conventional procedure described in the literature (e.g., Hemerik and Goeman, 2018), in which a randomization is performed by sampling a single transformation $G^{(b)}\sim\lambda$ and applying that to the entire sample $X_{1:n}$ (as a single multivariate observation). Our procedure is closer to that described by Romano (1989) and can be seen as a generalization of the conventional procedure to (possibly multivariate) i.i.d. observations. We recover the conventional procedure by treating $X_{1:n}$ as a single observation in the space $\mathbf{X}^{n}$ and taking the group action $G^{(b)}=G^{(b)}_{1}\times\dotsc\times G^{(b)}_{1}$ as an action on $\mathbf{X}^{n}$.
For the conventional procedure, an “unconditional” proof that the resulting test is level-$\alpha$ was given for discrete $\mathbf{G}$ by Hemerik and Goeman (2017). Their proof method can also be used without issue for (uncountable) compact $\mathbf{G}$. A second proof, in which the level of the test holds conditionally on the orbit of $X$, was given by Hemerik and Goeman (2018). That conditional proof, however, relies on the countability of the orbit, which fails when $\mathbf{G}$ is uncountable (e.g., the group of continuous rotations). Theorem 2.2 below fills this gap in the literature, holding for randomization tests of invariance for compact groups, conditioned on the orbit. Our proof relies on a sufficiency argument, thereby exposing randomization tests for invariance to be co-sufficient sampling methods (Barber and Janson, 2022). The high-level idea is as follows. A maximal invariant is a $\mathbf{G}$-invariant function, $M\colon\mathbf{X}\to\mathbf{M}$, that takes a different value on each orbit, so that $M(x)=M(x^{\prime})$ if and only if $gx=x^{\prime}$ for some $g\in\mathbf{G}$. It is known that any maximal invariant is a sufficient statistic for $\mathcal{P}^{\circ}(\mathbf{X})$, the set of $\mathbf{G}$-invariant probability distributions on $\mathbf{X}$ (Farrell, 1962; Dawid, 1985; Bloem-Reddy and Teh, 2020). Because a maximal invariant $M(X)$ uniquely identifies an orbit $O(X)$, the conditional distribution of $X$ given $M(X)$ is the distribution over $O(X)$ induced by pushing $\lambda$ (the uniform measure on $\mathbf{G}$) through the group action on that orbit. This holds for every invariant probability measure $P\in\mathcal{P}^{\circ}(\mathbf{X})$, even when the action of $\mathbf{G}$ on $\mathbf{X}$ has non-trivial stabilizers. Hence, conditioning on the observed maximal invariants and randomizing with $G\sim\lambda$ generates a set of conditionally i.i.d. samples, and the usual analysis of (conditional) Monte Carlo tests can be applied. Moreover, because every distribution in the null hypothesis set shares the same conditional distribution, $p_{B}$ is approximately uniformly distributed, with $p_{B}\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\to}}U\sim\textrm{Unif}[0,1]$ as $B\to\infty$.
Maximal invariants are not unique; we use $M$ to denote a generic maximal invariant, and $M(X)_{1:n}$ to denote the element-wise maximal invariants of $X_{1:n}$.
Theorem 2.2.
Let $\mathbf{G}$ be a compact group that acts on $\mathbf{X}$. Let $M$ be any maximal invariant. Assume that $X_{1:n}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}P$. For test statistic $T$ and randomization test in Algorithm 1, asume that $\mathbb{P}\left(T(X^{(b)}_{1:n})=T(X^{(b^{\prime})}_{1:n})\right)=0$ for $b\neq b^{\prime}$. For any fixed $B\in\mathbb{N}$ and $\alpha\in[0,1]$, if $P$ is $\mathbf{G}$-invariant then
[…]
The proof of Theorem 2.2 is in Section B.1. As noted by Dufour and Neves (2019), if $\alpha(B+1)$ is an integer, then the inequality in (4) becomes equality. If the probability of ties is non-zero (i.e., if $T$ is supported on a discrete set), then a modified version with randomized tie-breaking yields a level-$\alpha$ test (Dufour, 2006; Hemerik and Goeman, 2018).
@@
An important practical observation is that all of the above holds even for (possibly randomized) test statistics that depend on the observed data through $M(X)_{1:n}$. That means that functional parameters for $T$ can be selected in a data-driven way, as long as the procedure only uses $M(X)_{1:n}$. For example, the power of nonparametric tests based on kernel maximum mean discrepancy are known to depend greatly on the kernel function parameters (Ramdas et al., 2015), so our implementation relies on selecting parameters using the observed $M(X)_{1:n}$. This extends and further justifies an idea introduced by Biggs, Schrab and Gretton (2024) in the context of nonparametric permutation tests, which they allowed to depend on permutation-invariant functions of the data.
Example 1 (Test of 2D rotational invariance).
Consider a setting similar to the one visualized in the first column of Figure 1, where $X$ is a random vector in $\mathbb{R}^{2}$. Suppose that $X=ZR_{\theta}$, where $Z$ is a 2D Gaussian random variable centered on the point $(1,0)$ with identity covariance and $R_{\theta}$ is a random 2D rotation matrix of angle $\theta$,
$R_{\theta}=\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}\;.$ (5)
If $\theta$ is sampled over some range, it is straightforward to see that the marginal distribution of $X$ is invariant with respect to $\mathrm{SO}(2)$—the group of 2D rotations about the origin $(0,0)$—if and only if $\theta$ is uniform on $[0,2\pi)$.
Figure 2: Test rejection rates for the distribution described in Example 1 at sample size $n=100$ when $\theta$ is sampled uniformly from $[0,p)$. The distribution is $\mathrm{SO}(2)$-invariant only when $p=2\pi$. Test statistics (FS and SK) are described in Section 6; RN indicates randomization using Algorithm 1 and BS indicates a two-sample permutation test between $X_{1:n}$ and a single randomization generated as in Algorithm 1.
Using the kernel-based test statistics described in Section 6 within the randomization procedure in Algorithm 1, Figure 2 (RN-FS and RN-SK) shows the rejection rates, estimated over $1000$ simulations, of several tests for rotational invariance on samples $X_{1:n}$, where $\theta_{1},\dotsc,\theta_{n}$ are sampled i.i.d. over $[0,p)$ with $p\in\{\nicefrac{{\pi}}{{2}},\pi,\nicefrac{{3\pi}}{{2}},2\pi\}$. As $p\to 2\pi$, the distribution of $X$ becomes increasingly more invariant, which is reflected in the power of the tests decreasing with $p$. When $p=2\pi$, the tests reject at a rate of approximately $\alpha=0.05$. As a baseline, Figure 2 (BS-FS and BS-SK) also shows the results of a baseline test: comparing $X_{1:n}$ with a single randomization set $X_{1:n}^{(0)}$ using a two-sample permutation test. Histograms showing the $p$-values for these tests can be found in Section D.1.
3 Symmetry of conditional distributions
@@
$P_{Y|X}(B\mid x)=P_{Y|X}(gB\mid gx)\;,\quad x\in\mathbf{X},\ B\in\mathbf{S}_{\mathbf{Y}},\ g\in\mathbf{G}\;.$ (6)
@@
$P_{Y|X}(B\mid gx)=P_{Y|X}(B\mid x)\;,\quad x\in\mathbf{X},\ B\in\mathbf{S}_{\mathbf{Y}},\ g\in\mathbf{G}\;.$ (7)
As a simple example of equivariance, consider the case in which $Y\in\mathbb{R}^{d}$ is conditionally Gaussian-distributed and is acted on by $\text{SO}(d)$, the group of $d$-dimensional rotations.
@@
For random vectors $X$, $Y$ in $\mathbb{R}^{d}$, suppose that $Y|X\sim\mathsf{N}(X,\mathbf{I}_{d})$. It is straightforward to check that $P_{Y|X}$ is $\mathrm{SO}(d)$-equivariant. If $\varepsilon\sim\mathsf{N}(0,\mathbf{I}_{d})$ with $\varepsilon{\;\perp\!\!\!\perp\;}X$ so that $Y=X+\varepsilon$, then for any $g\in\mathrm{SO}(d)$, $A\subseteq\mathbb{R}^{d}$,
$P_{Y|X}(gY\in A\mid X)$ $=P_{Y|X}(gX+g\varepsilon\in A\mid X)$
$=P_{Y|X}(gX+\varepsilon\in A\mid X)=P_{Y|X}(Y\in A\mid gX)\;.$
If, instead, $Y|X\sim\mathsf{N}(\|X\|,\Sigma)$, then $P_{Y|X}$ is $\text{SO}(d)$-invariant.
We use the term “$\mathbf{G}$-symmetric” as a catch-all for either of the conditional symmetries in Definition 3.1. Similar to the hypotheses (1) in testing distributional invariance, the primary aim of this paper is to test the hypotheses
@@
3.1 Representatives and inversions
$[\mathbf{X}]$$(0,0)$$O(x)$$x=\tau(x)[x]$$[x]=\rho(x)$$\tau(x)$$M(x)$
Figure 3: A graphical depiction of a point $x\in\mathbb{R}^{2}$, its orbit $O(x)=\{x^{\prime}\in\mathbb{R}^{2}:\|x^{\prime}\|_{2}=M(x)\}$, and its orbit representative $[x]\in[\mathbf{X}]=\{(z,0):z\geq 0\}$, and the role of the orbit selector $\rho$, maximal invariant $M$, and representative inversion $\tau$ in the context of two-dimensional rotations about the origin $(0,0)$. Note that $[x]$ itself is also a maximal invariant.
One property distinguishing marginal invariance from conditional symmetry is that marginal invariance can only hold with respect to compact groups, whereas the group can be non-compact in the case of conditional symmetries. In principle, that makes conditional symmetries applicable to a wider range of problems, though it also makes them more challenging to formulate and analyze. In order to address those challenges, we rely on special structure arising from group actions, beyond what is typically needed for tests of distributional invariance. For technical details on measurability issues that can arise, the interested reader may refer to Section A.1. Throughout, we assume that the group action is proper so that conditions needed for measurability are satisfied.
On each orbit $O(x)$, we can choose an element $[x]\in O(x)$ as the orbit representative, so that $[x]=gx$ for some $g\in\mathbf{G}$. $[x]$ can be thought of as an “origin” on the orbit. The properties that we describe below do not depend on which element is chosen as the representative, but the underlying structure of the properties will be relative to that choice. For a particular choice of representatives, the subset of $\mathbf{X}$ consisting of each orbit’s representative is denoted by $[\mathbf{X}]$. Note that $[\mathbf{X}]\cap O(x)=\{[x]\}$. A function $\rho\colon\mathbf{X}\to[\mathbf{X}]$ that maps elements of $\mathbf{X}$ onto their corresponding orbit representatives is called an orbit selector. Any orbit selector is a maximal invariant by definition. Figures 3 and 4 depict these quantities for two- and three-dimensional rotations, respectively.
$O(x)$$[\mathbf{X}]$$(0,0,0)$$[x]=\rho(x)$$x$$g_{1}$$g_{2}$
Figure 4: A graphical depiction of a point $x\in\mathbb{R}^{3}$, its orbit $O(x)$ and orbit representative $[x]\in[\mathbf{X}]$ in the context of three-dimensional rotations about the origin $(0,0,0)$. Depicted are two rotations $g_{1}$ and $g_{2}$ that take $[x]$ to $x$. The inversion kernel $P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ is the uniform distribution over $\{g\in\mathrm{SO}(3):x=g[x]\}=g\mathbf{G}_{[x]}$.
@@ 3.1 Representatives and inversions
Recall that the stabilizer subgroup of $x$ is $\mathbf{G}_{x}=\{g\in\mathbf{G}:gx=x\}$, and that the action is free if $\mathbf{G}_{x}=\{\textrm{id}\}$ for all $x\in\mathbf{X}$. A $\mathbf{G}$-equivariant function $\tau\colon\mathbf{X}\to\mathbf{G}$ is called a representative inversion if $\tau(x)\rho(x)=x$ for all $x\in\mathbf{X}$. As the name suggests, $\tau$ returns the element of $\mathbf{G}$ that moves, or “inverts”, the representative $\rho(x)$ to $x$; the inverse group element $\tau(x)^{-1}$ moves $x$ to $\rho(x)$. In order for $\tau$ to be uniquely defined, the group action must be free. If it is not, an equivariant inversion probability kernel (inversion kernel for short) $P_{\tilde{\tau}|X}\colon\mathbf{X}\times\mathbf{S}_{\mathbf{G}}\to[0,1]$ can be used in place of $\tau$. If $\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$, then $x\stackrel{{\scriptstyle\text{\rm\tiny a.s.}}}{{=}}\tilde{\tau}\rho(x)$. At a high level, one may think of the inversion kernel $P_{\tilde{\tau}|X}$ as the uniform distribution on the left coset $g\mathbf{G}_{\rho(x)}$, where $g\rho(x)=x$. Assuming that the group action is proper (see Section A.1) ensures that the left coset is compact and a uniform distribution is well-defined. When the action is free, the inversion kernel simplifies to $P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)=\delta_{\tau(x)}({\,\vbox{\hbox{\tiny$\bullet$}}\,})$. Figures 3 and 4 depict these structures for $\text{SO}(2)$ and $\text{SO}(3)$ acting on $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$, respectively. The main difference is that the stabilizer subgroup of $x\in\mathbb{R}^{2}$ is trivial while that of $x\in\mathbb{R}^{3}$ is not.
In the setting of distributional invariance in Section 2.2, the conditional level stems from the decomposition $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}G\rho(X)$ with $G{\;\perp\!\!\!\perp\;}X$ and $G\sim\lambda$ (see Section B.1). To test for conditional symmetry, we will need a similar decomposition with a random element $\Gamma\in\mathbf{G}$, so that $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}\Gamma\rho(X)$. However, here the conditional distribution over $\mathbf{G}$ given an orbit of $X$ will in general depend on the orbit and not be invariant. We use $\Gamma$ to denote the random variable with this conditional distribution, which can be written as
$P_{\Gamma|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid[x])=\int_{\mathbf{X}}P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid x^{\prime})P_{X|[X]}(dx^{\prime}\mid[x])\;,$ (9)
where $P_{X|[X]}$ is the conditional distribution of $X$ given its orbit. We refer to $P_{\Gamma|[X]}$ as the group-orbit kernel as it is the distribution on $\mathbf{G}$ induced by restricting $P_{X}$ to the orbit $O([x])$. The notation indicates that it is conditioned on a particular orbit represented by $[x]$, but it can be induced by conditioning on any maximal invariant. The group-orbit kernel has the property that if $\Gamma\sim P_{\Gamma|[X]}$, then $(\rho(X),X)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\Gamma\rho(X))$. (See Lemma A.2 for a proof.)
In some cases, a representative inversion can still be (non-uniquely) defined when the action is not free (see Example 3), in which case an equivalent inversion kernel can be defined as $P_{\tilde{\tau}|X}^{\prime}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)\coloneqq P_{\tilde{\tau}|X}(\tau(x)^{-1}{\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid[x])$. Inversion kernels do not seem to have been used previously in statistics or machine learning, likely owing to their relatively recent appearance in probability (Kallenberg, 2011). However, the deterministic representative inversions have been seen in the recent machine learning literature (Bloem-Reddy and Teh, 2020; Winter et al., 2022; Kaba et al., 2023).
As previously mentioned, when $\mathbf{G}$ is non-compact, there may not exist measurable maximal invariants, orbit selectors, and inversion kernels. We assume that the action of $\mathbf{G}$ on any set is continuous and proper; these conditions are implicit in statements such as “let $\mathbf{G}$ be a group that acts on $\mathbf{X}$”. For our purposes, properness is sufficient for the existence of measurable versions of these objects. See Section A.1 for more details.
@@
The following examples illustrate the main ideas, and provide background for our experiments in Section 7.

### v2

@@ 2.2 Testing for distributional invariance
The literature on hypothesis tests for group invariance of a probability measure is extensive, going back at least to Hoeffding (1952), and even earlier for special cases. A textbook treatment of such tests when $\mathbf{G}$ is finite can be found in (Lehmann and Romano, 2005, Ch. 15.2); although the technical details for infinite compact groups are more involved, the basic structure is similar. As a reference point for the methods we will develop for testing conditional symmetry, we briefly review the basics of testing for distributional invariance, highlighting particular aspects that are important in the sequel.
A function $f$ with domain $\mathbf{X}$ is $\mathbf{G}$-invariant if $f(gx)=f(x)$, all $x\in\mathbf{X},g\in\mathbf{G}$. Observe that a $\mathbf{G}$-invariant function is constant on each orbit. This specializes to probability measures.
@@
A probability measure $P$ on $\mathbf{X}$ is $\mathbf{G}$-invariant if $g_{*}P(A)=P(A)$ for all $g\in\mathbf{G}$, $A\in\mathbf{S}_{\mathbf{X}}$.
When the group action is continuous, $\mathbf{G}$-invariance of a probability measure is possible only when $\mathbf{G}$ is compact (a non-compact group acting continuously would induce non-compact orbits, which would not admit invariant probability measures). For a specified compact group $\mathbf{G}$, consider the hypotheses
$H_{0}\colon P\text{ is $\mathbf{G}$-invariant}\quad\text{versus}\quad H_{1}\colon P\text{ is not $\mathbf{G}$-invariant.}$ (1)
The standard level-$\alpha$ test based on statistic $T$ is
$\phi_{\alpha}(X)=\mathbbm{1}\{T(X)>q_{T(GX)}(1-\alpha)\}\;,$
where $q_{T(GX)}(1-\alpha)$ is the quantile function of $T(GX)$ at $1-\alpha$, with $X\sim P$ and $G\sim\lambda$, a random element of $\mathbf{G}$ sampled from Haar measure (i.e., the uniform distribution on $\mathbf{G}$). The usual way to show that this test has level $\alpha$ uses the fact that $P$ is $\mathbf{G}$-invariant if and only if $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}GX$. Hence $\mathbb{E}[\phi_{\alpha}(X)]=\mathbb{E}[\phi_{\alpha}(GX)]\leq\alpha$. For small and finite $\mathbf{G}$, the quantile $q_{T(GX)}(1-\alpha)$ can be computed exactly by applying each element of $\mathbf{G}$ to $X$. However, in the case that $\mathbf{G}$ is large or uncountable, randomization tests (also called “stochastic approximation tests,” Lehmann and Romano, 2005) may be used to obtain a Monte Carlo estimate of the quantile. Algorithm 1 shows the basic recipe when the elements of $X_{1:n}$ are i.i.d. [^2] If, instead, $X_{1:n}$ is treated as a single observation of a multivariate random variable (e.g., Hemerik and Goeman, 2018), then randomization is performed by applying a single random transformation $G^{(b)}\sim\lambda$ to the entire sample $X_{1:n}$. In the basic scheme, a fixed number $B$ of random group elements are sampled independently from Haar measure. As with all randomization tests, in order to avoid potential issues related to increasing the Type I error rate due to the randomness of the test, $B$ should be set sufficiently large. When sampling is computationally expensive, there are methods for selecting $B$ adaptively in order to reduce computation while maintaining statistical validity (see Stoepker and Castro, 2024, for a recent review), and recent work has investigated more statistically efficient sampling schemes for permutation tests (Ramdas et al., 2023a; Koning and Hemerik, 2024).
@@
$p_{B}\coloneqq\frac{1+\sum_{b=1}^{B}\mathbbm{1}\{T(X_{1:n}^{(b)})\geq T(X_{1:n})\}}{1+B}$
@@
Somewhat less conventionally, it can be shown that Algorithm 1 instantiates a test that is level-$\alpha$ conditionally on $M(X)_{1:n}$, the element-wise maximal invariants of $X_{1:n}$, which is a sufficient statistic for the set of $\mathbf{G}$-invariant probability measures (Farrell, 1962; Dawid, 1985; Bloem-Reddy and Teh, 2020). Hence, if $X_{1:n}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}P$ and $P$ is $\mathbf{G}$-invariant, then for any $\alpha\in[0,1]$,
$\mathbb{P}\left(p_{B}\leq\alpha\mid M(X)_{1:n}\right)=\frac{\lfloor\alpha(B+1)\rfloor}{B+1}\leq\alpha\;,\quad P\text{-a.s.}$ (2)
An orbit selector is especially useful in this setting. With an orbit selector $\rho$ fixed, it is known that every $\mathbf{G}$-invariant probability measure factorizes via the distributional identity $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}G\rho(X)$, with $G{\;\perp\!\!\!\perp\;}X$ and $G\sim\lambda$ (see Appendix B). Once the orbit is known, all remaining variation in $X$ is produced by $G\sim\lambda$. Hence, the sufficiency of $\rho(X)$ and the simplicity of Algorithm 1.
@@
Conditioning the test on $\rho(X)_{1:n}$ has important practical implications for increasing test power. Because (2) holds conditioned on $\rho(X)_{1:n}$, the statistic $T$ may depend on the observed data through $\rho(X)_{1:n}$, meaning that functional parameters of $T$ may be selected in a data-adaptive way, as long as the selection only uses $\rho(X)_{1:n}$ (or any other $\mathbf{G}$-invariant function of $X_{1:n}$). For example, the power of nonparametric tests based on kernel maximum mean discrepancy is known to depend greatly on the kernel function parameters such as bandwidth (Ramdas et al., 2015), so power may improve when parameters are selected using the observed maximal invariants. This provides a complementary justification for an idea introduced by Biggs, Schrab and Gretton (2024) in the context of nonparametric permutation tests, which they allowed to depend on permutation-invariant functions of the data. We will build on these ideas when constructing tests for conditional symmetry.
Although (2) has been proven for special cases (Romano, 1989; Hemerik and Goeman, 2018), the literature seems to be missing the sufficiency perspective, as well as a full proof of the level of the conditional test for a compact group that may not act freely (e.g., the group of continuous rotations for $\mathbb{R}^{d}$, $d\geq 3$), though Romano (1989, Example 2) outlines such an argument for rotations. For completeness, we provide a formal statement and proof in Appendix B.
2.3 Symmetry of conditional distributions
@@
$P_{Y|X}(B\mid x)=P_{Y|X}(gB\mid gx)\;,\quad x\in\mathbf{X},\ B\in\mathbf{S}_{\mathbf{Y}},\ g\in\mathbf{G}\;.$
@@
$P_{Y|X}(B\mid gx)=P_{Y|X}(B\mid x)\;,\quad x\in\mathbf{X},\ B\in\mathbf{S}_{\mathbf{Y}},\ g\in\mathbf{G}\;.$
We use the term $\mathbf{G}$-symmetric” as a catch-all for either of the conditional symmetries in Definition 2.2. As a simple example of equivariance, consider a case in which $Y\in\mathbb{R}^{d}$ is conditionally Gaussian-distributed and is acted on by $\mathbb{T}_{d}$, the locally compact group of $d$-dimensional translations.
@@
For random vectors $X$, $Y$ in $\mathbb{R}^{d}$, suppose that $Y|X\sim\mathsf{N}(X,\Sigma)$, where $\Sigma$ is some fixed covariance matrix. It is straightforward to check that $P_{Y|X}$ is $\mathbb{T}_{d}$-equivariant. If $\varepsilon\sim\mathsf{N}(0,\Sigma)$ with $\varepsilon{\;\perp\!\!\!\perp\;}X$ so that $Y\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}X+\varepsilon$, then for any $g_{v}\in\mathbb{T}_{d}$ where $g_{v}x\coloneqq v+x$ for $v\in\mathbb{R}^{d}$, $A\subseteq\mathbb{R}^{d}$,
$P_{Y|X}(g_{v}Y\in A\mid X)=P_{Y|X}(v+X+\varepsilon\in A\mid X)=P_{Y|X}(Y\in A\mid g_{v}X)\;.$
If, instead, $Y|X\sim\mathsf{N}(1_{d},\Sigma)$, where $1_{d}$ is the $d$-dimensional vector of ones, then $P_{Y|X}$ is $\mathbb{T}_{d}$-invariant. In this example, conditional invariance is equivalent to independence. That is not a coincidence, but as we establish in Theorem 3.1, is a consequence of the transitivity of $\mathbb{T}_{d}$ acting on $\mathbf{X}=\mathbb{R}^{d}$.
Similar to the hypotheses (1) in testing distributional invariance, the primary aim of this paper is to test the hypotheses
@@
In the setting of distributional invariance from Section 2.2, the test properties stem from the decomposition $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}G\rho(X)$ with $G{\;\perp\!\!\!\perp\;}X$ and $G\sim\lambda$. To test for conditional symmetry in a similar fashion, we require an analogous decomposition of the joint distribution of $(X,Y)$ in terms of a random element $\Gamma\in\mathbf{G}$, to be shared between $X$ and $Y$. The conditional setting presents new challenges, however. Conditional symmetries can hold with respect to non-compact groups, as in Example 1. Hence, in contrast to tests for distributional invariance, tests for conditional symmetry cannot simply rely on the sampling from a normalized Haar measure, which does not exist for non-compact groups. Moreover, because equivariance encodes systematic regularities in the relationship between variables, the required random group element $\Gamma$ cannot be independent from $X$ and $Y$. In short, whereas the test for distributional invariance required only a maximal invariant and a uniformly random group element, tests for conditional symmetry must make use of additional structure arising from the group actions.
3 Characterizations of conditional symmetry
In addition to orbit selectors, the tests we will develop rely on a technique known as group inversion. [^3] For technical details on measurability issues that can arise, the interested reader may refer to Section A.1. In particular, measurable versions of the objects defined below may not exist. Throughout, we assume that the group action is proper so that conditions needed for measurability are satisfied. Recall from Section 2.2 that an orbit selector $\rho$ maps an element $x$ to its orbit representative, $\rho(x)=r_{x}$. This is well-defined even for non-compact groups, subject to the regularity condition of proper action. The properties that we describe below do not depend on which elements are chosen as the representatives, but the specific subsequent functions and kernels will be relative to that choice.
@@ 3 Characterizations of conditional symmetry
Recall that the stabilizer subgroup of $x$ is $\mathbf{G}_{x}=\{g\in\mathbf{G}:gx=x\}$, and that the action is free if $\mathbf{G}_{x}=\{\textrm{id}\}$ for all $x\in\mathbf{X}$. A $\mathbf{G}$-equivariant function ${\psi^{\circ}}\colon\mathbf{X}\to\mathbf{G}$ is called a representative inversion if ${\psi^{\circ}}(x)\rho(x)=x$ for all $x\in\mathbf{X}$. As the name suggests, $\psi^{\circ}$} returns the element of $\mathbf{G}$ that moves, or “inverts”, the representative $\rho(x)$ to $x$; the inverse group element ${\psi^{\circ}}(x)^{-1}$ moves $x$ to $\rho(x)$. In order for $\psi^{\circ}$} to be uniquely defined, the group action must be free. If it is not, an equivariant inversion probability kernel (inversion kernel for short) $P_{\Psi|X}\colon\mathbf{X}\times\mathbf{S}_{\mathbf{G}}\to[0,1]$ can be used in place of $\psi^{\circ}$}. If $\Psi|X\sim P_{\Psi|X}$, then $X\stackrel{{\scriptstyle\text{\rm\tiny a.s.}}}{{=}}\Psi\rho(X)$. At a high level, one may think of the inversion kernel $P_{\Psi|X}$ as the uniform distribution on the left coset $g\mathbf{G}_{\rho(X)}$, where $g\rho(X)=X$. Assuming that the group action is proper (see Section A.1) ensures that the left coset is compact and a uniform distribution is well-defined. When the action is free, the inversion kernel simplifies to $P_{\Psi|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)=\delta_{{\psi^{\circ}}(x)}({\,\vbox{\hbox{\tiny$\bullet$}}\,})$. Figure 2 visualizes these structures for $\mathrm{SO}(2)$ and $\mathrm{SO}(3)$. The main difference is that the stabilizer subgroup of $x\in\mathbb{R}^{2}$ is trivial while that of $x\in\mathbb{R}^{3}$ is not. In some cases, a representative inversion can still be (non-uniquely) defined when the action is not free (see Example 2), in which case an equivalent inversion kernel can be defined as $P_{\Psi|X}^{\prime}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)\coloneqq P_{\Psi|X}({\psi^{\circ}}(x)^{-1}{\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid\rho(x))$.
@@
The following examples illustrate the main ideas and provide background for our experiments in Section 6.